# Function-Space Learning Rates

### v1

@@ 2 Related work
Perhaps the earliest and best-known work on hyperparameter scaling is µP (Yang & Hu, 2022). µP derives how to scale random initialisations and learning rates as you increase model width, such that the function-space learning rates remain asymptotically constant (i.e. the magnitude of the activations does not blow up to infinity or shrink to zero as the model gets wider). µP has since been extended to depth-scaling (Yang et al., 2023; Bordelon et al., 2023) and networks trained with sharpness-aware minimisation (Haas et al., 2024), and is closely related to the mean-field analysis of neural networks grounded in statistical mechanics (Mei et al., 2018; Geiger et al., 2020; Bordelon & Pehlevan, 2022; Rotskoff & Vanden‐Eijnden, 2022). Because this approach derives the function-space learning rates analytically, (in contrast to our approach of measuring them), it requires restrictive assumptions, including that the network is wide, and close to a random initialisation. Extending these results to more general cases is made complicated by the fact that they typically rely on heavy-duty mathematical machinery such as Tensor-Programs (Yang, 2021b, 2020, a; Yang & Hu, 2022; Yang et al., 2022, 2023) or dynamical mean-field theory (Bordelon & Pehlevan, 2022, 2023; Bordelon et al., 2023, 2024)
@@
Finally, Everett et al. (2024) recently evaluated the “alignment” assumptions (concerning the size of dot products between activations and updates across different layers) made by µP and mean-field parametrisations (Yang & Hu, 2022; Mei et al., 2018; Bordelon & Pehlevan, 2022), and found that they are often violated in practice. They showed empirically that alignment in real models is highly dynamic and complex throughout training, meaning that static parametrisations like µP and mean-field could be problematic. By contrast, we propose methods that can directly measure the function-space learning rates throughout training, avoiding the need for such assumptions.
@@ 3.1 Monte-Carlo estimation of the layerwise function-space learning rate
At the core of our contributions is the estimation of the layerwise function-space learning rates, i.e. the magnitude of the change in output logits arising from a particular change in the $\ell^{\text{th}}$ parameter tensor. We begin by considering the full change in the function output, $f_{nk}$. Here, $n$ indexes the $N$ datapoints in the minibatch, and $k$ indexes the $K$ output features. We use $\Delta_{\ell}f_{nk}$ to denote a first-order Taylor approximation of the change in the outputs due to a particular change, $\Delta\mathbf{W}_{\ell}$, in the $\ell$th parameter, $\mathbf{W}_{\ell}$,
@@
Here, $W^{\ell}_{ij}$ is the ${ij}$th element of $\mathbf{W}_{\ell}$ and $\frac{df_{nk}}{dW^{\ell}_{ij}}$ is the gradient of the output w.r.t. $W^{\ell}_{ij}$. Note that for ease of notation we assume $\mathbf{W}_{\ell}$ is a matrix, but $\mathbf{W}_{\ell}$ can be a tensor of any rank. We are interested in the layerwise function-space learning rate, i.e. the RMS norm of $\Delta_{\ell}\mathbf{f}$,
@@
        $\eta^{\ell}\leftarrow\eta_{0}\left\lVert\Delta_{\ell}\mathbf{f}\right\rVert_{\text{RMS}}^{(t)}/\left\lVert\Delta_{\ell}\mathbf{f}\right\rVert_{\text{RMS}}^{\text{(base,$t$)}}$
@@ 3.2 More efficient function-space learning rate estimates using Kronecker factorisation
The approach in Sec. 3.1 requires one backward pass per sample, which could still be inefficient if we need multiple samples for a good estimate. We remedy this by exploiting the structure of $\phi$. Noting that $i,j$ index the rows and columns of $\mathbf{W}_{\ell}$, we rewrite Eq. (4) as,
@@
Instead, we assume a Kronecker-factored covariance matrix (Martens & Grosse, 2015). Specifically, we assume that we have a covariance matrix $\mathbf{U}$ over rows and $\mathbf{V}$ over columns, giving
@@
$\left\lVert\Delta_{\ell}\mathbf{f}\right\rVert_{\text{RMS}}^{2}$ $=$ $=\frac{\mathbb{E}\left(\sum_{ii}[\mathbf{Z}\mathbf{Z}^{T}]_{ii^{\prime}}\right)\mathbb{E}\left(\sum_{jj^{\prime}}[\mathbf{Z}^{T}\mathbf{Z}]_{jj^{\prime}}\right)}{\mathbb{E}[||\mathbf{Z}||_{\mathcal{F}}^{2}]}$ (16)
@@ 3.3 Function-space learning rate matching (FLeRM)
FLeRM uses the machinery developed above to record the function-space learning rates in a small, cheap base model. Then, in the larger, more expensive model, FLeRM sets the parameter-space learning rates such that the function-space learning rates match those in the base model (see Alg. 1, blue text, for details).
@@ 4 Experiments
Full details of the models [^1] We provide our code at https://github.com/edwardmilsom/function-space-learning-rates-paper used in the experiments can be found in Appendix A. The base ResMLP is an MLP with 4 hidden layers, each with residual connections, trained for 50 epochs on flattened CIFAR10 images (Krizhevsky & Hinton, 2009). The base transformer is decoder-only, has two self-attention + feedforward blocks (Vaswani et al., 2017), and is trained on a subset of the Wikitext-103 dataset (Merity et al., 2016). We compare 3 different types of Layernorm (Ba et al., 2016) in the transformers, with affine transformations disabled: PostNorm is $\text{Norm}(x+f(x))$ , PreNorm is $x+f(\text{Norm}(x)$, and PreNormPostMod is $x+\text{Norm}(f(x))$. In the ResMLP, scaling width increases the hidden dimension, whilst in the transformers, we scale the embedding dimension, the feedforward hidden dimension, and the number of heads. In all models, depth scaling increases the number of “residual blocks” that form the hidden layers of the model. Both ResMLP and the transformers used the Adam optimiser (Kingma, 2014) with a constant learning rate schedule. In all plots, train loss is averaged over the last 200 batches of training.
@@ 4.1 Analysing function-space learning rates
In Figure 1 we measure the function-space learning rates using the techniques presented in Sections 3.1 and 3.2 for the ResMLP and PostNorm transformer models. Plots for the PreNorm and PreNormPostMod transformer can be found in Figure 6, though they are very similar. We use 40 batches of data to warm up the EMAs as suggested Section 3.2, then measure the function-space learning rate every 100 iterations. We use an EMA decay rate of $\beta=0.9$. that in both models, the function-space learning rates change over time, despite the parameter-space learning rates being fixed. The most obvious pattern is that the function space learning rates fall monotonically for all parameters, except the input embedding layer, revealing an implicit scheduling of the function-space learning rates under standard Adam training. Interestingly, in the ResMLP, whilst the function-space learning rates initially decay, the hidden layers and input layer eventually start increasing again, whilst the output layer plateaus. In the transformer, the embedding layer’s function-space learning rate actually increases over time. One possible explanation is that the noisy initialisation in hidden and output layers effectively scrambles the signal from the input layer, but as these layers are trained, the input embedding can have a clearer, stronger effect on the output.
@@ A.1 ResMLP
We optimise the model using Adam with default settings (other than the learning rate which we set using the methods detailed in this paper). We train for 50 epochs.

### v2

@@ 2 Related work
Perhaps the earliest and best-known work on hyperparameter scaling is µP (Yang & Hu, 2022). µP derives how to scale random initialisations and learning rates as you increase model width, such that the function-space learning rates remain asymptotically constant (i.e. the magnitude of the activations does not blow up to infinity or shrink to zero as the model gets wider). µP has since been extended to depth-scaling (Yang et al., 2023; Bordelon et al., 2023) and networks trained with sharpness-aware minimisation (Haas et al., 2024), and is closely related to the mean-field analysis of neural networks grounded in statistical mechanics (Mei et al., 2018; Rotskoff & Vanden‐Eijnden, 2022; Sirignano & Spiliopoulos, 2019; Chizat & Bach, 2018; Geiger et al., 2020; Bordelon & Pehlevan, 2022). Because this approach derives the function-space learning rates analytically, (in contrast to our approach of measuring them), it requires restrictive assumptions, including that the network is wide, and close to a random initialisation. Extending these results to more general cases is made complicated by the fact that they typically rely on heavy-duty mathematical machinery such as Tensor-Programs (Yang, 2021b, 2020, a; Yang & Hu, 2022; Yang et al., 2022, 2023) or dynamical mean-field theory (Bordelon & Pehlevan, 2022, 2023; Bordelon et al., 2023, 2024).
@@
Chizat & Netrapalli (2024) quantify feature learning in neural networks as the angle between feature updates and backward passes, enabling analysis of hyperparameter scaling laws and development of improved scaling rules for deep networks. By contrast, our work directly measures the function-space learning rates using autodiff and a Monte-Carlo approximation.
Finally, Everett et al. (2024) recently showed empirically that alignment (concerning the size of dot products between activations and updates across different layers) in real models is highly dynamic and complex throughout training. This can make choosing the correct alignment assumptions in µP and mean-field parametrisations (Yang & Hu, 2022; Mei et al., 2018; Bordelon & Pehlevan, 2022), a very difficult task. By contrast, we propose methods that can directly measure the function-space learning rates throughout training, avoiding the need for such assumptions and analysis.
@@ 3.1 Monte-Carlo estimation of the layerwise function-space learning rate
At the core of our contributions is the estimation of the layerwise function-space learning rates, i.e. the magnitude of the change in output logits arising from a particular change in the $\ell^{\text{th}}$ parameter tensor. We begin by considering the full change in the function output, $f_{nk}$. Here, $n$ indexes the $N$ datapoints in the minibatch, and $k$ indexes the $K$ output features. We use $\Delta_{\ell}f_{nk}$ to denote a first-order Taylor approximation of the change in the outputs due to a particular change, $\Delta\mathbf{W}^{\ell}$, in the $\ell$th parameter, $\mathbf{W}^{\ell}\in\mathbb{R}^{N_{\ell}\times N_{\ell-1}}$,
@@
Here, $W^{\ell}_{ij}$ is the ${ij}$th element of $\mathbf{W}^{\ell}$ and $\frac{df_{nk}}{dW^{\ell}_{ij}}$ is the gradient of the output w.r.t. $W^{\ell}_{ij}$. Note that for ease of notation we assume $\mathbf{W}^{\ell}$ is a matrix, but $\mathbf{W}^{\ell}$ can be a tensor of any rank. We are interested in the layerwise function-space learning rate, i.e. the RMS norm of $\Delta_{\ell}\mathbf{f}$,
@@
        $\eta^{\ell}\leftarrow\eta_{0}\left\lVert\Delta_{\ell}\mathbf{f}\right\rVert_{\text{RMS}}^{(\text{base},t)}/\left\lVert\Delta_{\ell}\mathbf{f}\right\rVert_{\text{RMS}}^{\text{($t$)}}$
@@ 3.2 More efficient function-space learning rate estimates using Kronecker factorisation
The approach in Sec. 3.1 requires one backward pass per sample, which could still be inefficient if we need multiple samples for a good estimate. We remedy this by exploiting the structure of $\phi$. Noting that $i,j$ index the rows and columns of $\mathbf{W}^{\ell}$, we rewrite Eq. (4) using $\mathbf{Z}\in\mathbb{R}^{N_{\ell}\times N_{\ell-1}}$,
@@
Instead, we assume a Kronecker-factored covariance matrix (Martens & Grosse, 2015). Specifically, we assume that we have a covariance matrix $\mathbf{U}\in\mathbb{R}^{N_{\ell}\times N_{\ell}}$ over rows and $\mathbf{V}\in\mathbb{R}^{N_{\ell-1}\times N_{\ell-1}}$ over columns, giving
@@
$\left\lVert\Delta_{\ell}\mathbf{f}\right\rVert_{\text{RMS}}^{2}$ $=$ $=\frac{\mathbb{E}\left(\sum_{ii^{\prime}}[\mathbf{Z}\mathbf{Z}^{T}]_{ii^{\prime}}\right)\mathbb{E}\left(\sum_{jj^{\prime}}[\mathbf{Z}^{T}\mathbf{Z}]_{jj^{\prime}}\right)}{\mathbb{E}[||\mathbf{Z}||_{\mathcal{F}}^{2}]}$ (16)
@@ 3.3 Function-space learning rate matching (FLeRM)
FLeRM uses the machinery developed above to record the function-space learning rates $\left\lVert\Delta_{\ell}\mathbf{f}\right\rVert_{\text{RMS}}^{(\text{base},t)}$ at iteration $t$ in a small, cheap base model. Then, in the larger, more expensive model, FLeRM uses the ratio between the current function-space learning rates $\left\lVert\Delta_{\ell}\mathbf{f}\right\rVert_{\text{RMS}}^{(t)}$ and the base model function-space learning rates $\left\lVert\Delta_{\ell}\mathbf{f}\right\rVert_{\text{RMS}}^{(\text{base},t)}$ to set the parameter-space learning rates at time $t$, such that the function-space learning rates match those in the base model (see Alg. 1, blue text, for details). Note that for efficiency reasons, this is usually done at regular intervals or once at the start of training, rather than at every iteration.
@@ 4 Experiments
Full details of the models [^1] We provide our code at https://github.com/edwardmilsom/function-space-learning-rates-paper used in the experiments can be found in Appendix A. The base ResMLP is an MLP with 4 hidden layers, each with residual connections, trained for 50 epochs on flattened CIFAR-10 images (Krizhevsky & Hinton, 2009). The base transformer is decoder-only, has two self-attention + feedforward blocks (Vaswani et al., 2017), and is trained on a subset of the Wikitext-103 dataset (Merity et al., 2016). The widest transformer has roughly 814M parameters. We compare 3 different types of Layernorm (Ba et al., 2016) in the transformers, with affine transformations disabled: PostNorm is $\text{Norm}(x+f(x))$ , PreNorm is $x+f(\text{Norm}(x))$, and PreNormPostMod is $x+\text{Norm}(f(x))$. In the ResMLP, scaling width increases the hidden dimension, whilst in the transformers, we scale the embedding dimension, the feedforward hidden dimension, and the number of heads. In all models, depth scaling increases the number of “residual blocks” that form the hidden layers of the model. Both ResMLP and the transformers used the Adam optimiser (Kingma, 2014) with a constant learning rate schedule. In all plots, train loss is averaged over the last 200 batches of training.
@@ 4.1 Analysing function-space learning rates
In Figure 1 we measure the function-space learning rates using the techniques presented in Sections 3.1 and 3.2 for the ResMLP and PostNorm transformer models. Plots for the PreNorm and PreNormPostMod transformer can be found in Figure 6, though they are very similar. We use 40 batches of data to warm up the EMAs as suggested in Section 3.2, then measure the function-space learning rate every 100 iterations. We use an EMA decay rate of $\beta=0.9$.
In Figure 1 we see that in both models, the function-space learning rates change over time, despite the parameter-space learning rates being fixed. The most obvious pattern is that the function space learning rates fall monotonically for all parameters, except the input embedding layer, revealing an implicit scheduling of the function-space learning rates under standard Adam training. Interestingly, in the ResMLP, whilst the function-space learning rates initially decay, the hidden layers and input layer eventually start increasing again, whilst the output layer plateaus. In the transformer, the embedding layer’s function-space learning rate actually increases over time. One possible explanation is that the noisy initialisation in hidden and output layers effectively scrambles the signal from the input layer, but as these layers are trained, the input embedding can have a clearer, stronger effect on the output.
@@ 5 Conclusion
Acknowledgements
Edward Milsom and Ben Anson are funded by the Engineering and Physical Sciences Research Council via the COMPASS Centre for Doctoral Training at the University of Bristol. This work was carried out using the computational facilities of the Advanced Computing Research Centre, University of Bristol - http://www.bris.ac.uk/acrc/. We would like to thank Dr. Stewart for GPU compute resources.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ A.1 ResMLP
We optimise the model using Adam with default settings (other than the learning rate which we set using the methods detailed in this paper). We train for 50 epochs on the CIFAR-10 dataset (Krizhevsky & Hinton, 2009).
@@ C.5 LoRA
C.6 Bias and variance of estimator under difference covariance assumptions
In Figure 15 we show that our proposed Kronecker-factored assumption in Section 3.2 greatly reduces the variance of the function-space learning rate estimator whilst avoiding the large bias an oversimplifying assumption such as an IID covariance structure would introduce. Here the “no assumption” estimator is used as the “true” value for computing the biases, as it is an unbiased estimator.
Figure 15: Comparison of bias and variance of the function-space learning rate estimator when using different covariance matrix / dependence assumptions in Section 3.1 / 3.2. Computed over 10,000 batches to estimate the FSLRs at the first step of training for a feedforward layer in the network. Bias is taken as the absolute difference from “No assumption”, since no assumption (i.e. ignoring Section 3.2 and estimating the quantity directly) is an unbiased estimator. We see that an IID assumption has extremely low variance (not even visible on the plot), but very high bias, whilst the unbiased “no assumption” estimator has very high variance. The KFAC estimator proposed in Section 3.2 has a small amount of bias and much smaller variance than the “no assumption” estimator.
C.7 Test loss
In Figure 16 we show the width transfer plot from the main text but with test losses instead of train losses. As expected, the test loss plots for the Transformer models look very similar to the main text, since we are only training for 1 epoch and therefore expect test loss to look like train loss. However, the ResMLP is trained for multiple epochs on CIFAR-10, and so complex overfitting patterns occur, which probably explains why the test loss curves do not line up perfectly, though the optima are closer together with FLeRM than with standard practice.
Figure 16: Width transfer plot for plotting test loss instead of train loss.
C.8 Cosine Annealing Scheduler
For simplicity we used a constant LR scheduler in the main text. To verify that our method works with LR schedulers, we reran the width transfer experiments using a cosine annealing scheduler. Since FLeRM modifies the layerwise learning rates, we record the ratio of the current scheduled learning rate with the starting learning rate in the base model, then apply these ratios as the scheduler in the scaled model. This is equivalent to scheduling the ”target” function-space learning rate for each layer. The results are shown in Figure 17.
Figure 17: Width transfer plot for Transformer (PreNormPostMod) using a CosineAnnealingLR scheduler.
C.9 Elementwise Affine Transformations
For simplicity in prototyping, elementwise affine transformations in layernorms were disabled in the main experiments. To verify our method still works with these enabled, we reran the width transfer experiment for the PreNormPostMod transformer, and, as seen in Figure 18, found no issues.
Figure 18: Width transfer plot for Transformer (PreNormPostMod) with elementwise affine transformations enabled in the Layernorms.
C.10 Simultaneous Width + Depth Scaling
To test whether our method still works when simultaneously scaling width and depth, we ran the PreNormPostMod transformer experiment again, scaling the width+depth, and found that hyperparameter transfer still holds in this setting. The results are shown in Figure 19
Figure 19: Width + Depth simultaneous scaling transfer plot for Transformer (PreNormPostMod).
C.11 Other optimisers
To test whether our method still works when using optimisers with Adam, we repeated the Transformer (PreNormPostMod) width transfer experiments for a range of different optimisers.
In Figure 20 we used SGD with momentum. In Figure 21 we used SignSGD. In Figure 22 we used AdamW. In Figure 23, we used Adamax. In Figure 24, we used Adagrad. We can see that FLeRM improves hyperparameter transfer in all cases. In SGD, the curves are very noisy both with and without FLeRM, matching the common practice of not using SGD for transformers. However, even though the curves are less neat than the other optimisers, we still see that FLeRM aligns the curves better.
To demonstrate this is a problem with using SGD with transformers and not FLeRM, in Figure 25 we ran the ResMLP width transfer experiment with SGD + momentum, where we observely that FLeRM very cleanly improves hyperparameter transfer.
Figure 20: Width transfer plot for Transformer (PreNormPostMod) using SGD instead of Adam. Momentum 0.9.
Figure 21: Width transfer plot for Transformer (PreNormPostMod) using SignSGD / Signum instead of Adam. Momentum 0.9.
Figure 22: Width transfer plot for Transformer (PreNormPostMod) using AdamW instead of Adam. Weight decay 0.1.
Figure 23: Width transfer plot for Transformer (PreNormPostMod) using Adamax instead of Adam.
Figure 24: Width transfer plot for Transformer (PreNormPostMod) using Adagrad instead of Adam.
Figure 25: Width transfer plot for ResMLP using SGD instead of Adam. Momentum 0.9.
Appendix D Exploiting known structure for lower variance function-space learning rate estimation
A typical neural network ends in a “readout” layer $\{\mathbf{W}^{L},\mathbf{b}^{L}\}$, mapping from the hidden dimension $d_{\text{model}}$ to the number of output classes / vocab size $K$ via the transformation
$\mathbf{f}_{n}=\mathbf{W}^{L}\mathbf{h}^{L-1}_{n}+\mathbf{b}^{L}$ (29)
where $\mathbf{h}^{L-1}_{n}$ is the output of the previous layer for the $n^{\text{th}}$ datapoint.
This layer can be very large if $K$ is large (the vocab size of an LLM), and can have large effects on the training dynamics of the neural network (Fig. 1). It is therefore important to ensure its function-space learning rate estimate is as accurate as possible. Luckily, it turns out that by nature of being the final layer, there is extra independence structure in the function-space learning rate estimation procedure we can exploit to lower the variance.
Subbing Eq. 29 into Eq. 3, we have
$\phi=\frac{1}{\sqrt{NK}}\sum_{nk}\omega_{nk}\left(b_{k}+\sum_{\alpha}W_{k\alpha}^{L}h_{n\alpha}^{L-1}\right)$ (30)
and subsequently subbing Eq. 3 into Eq. 8, considering the function-space learning rate for the readout weight matrix $\mathbf{W}^{L}\in\mathbb{R}^{K\times d_{\text{model}}}$, we obtain
$Z_{ij}^{(\mathbf{W}^{L})}=\Delta W_{ij}^{L}\frac{1}{\sqrt{NK}}\sum_{n}\omega_{ni}h^{L-1}_{ni}.$ (31)
Note that this only includes a single random variable $\omega_{ni}$. This means the random variables $\{Z_{ij}^{(\mathbf{W}^{L})}\}_{ij}$ are independent for different values of $i$ (i.e. between different rows), but dependent for different values for $j$ (i.e. between different columns). Hence we can assume independence between the rows of $Z_{ij}^{(\mathbf{W}^{L})}$, drastically reducing the variance of our estimator.
In particular, this corresponds to assuming that $\mathbf{U}$ is diagonal in Section 3.2, which results in $||\Delta_{\ell}\mathbf{f}||^{2}_{\text{RMS}}=\left(\sum_{i}U_{ii}\right)\left(\sum_{jj^{\prime}}V_{jj^{\prime}}\right)=\text{tr}(\mathbf{U})\left(\sum_{jj^{\prime}}V_{jj^{\prime}}\right)$ and therefore our estimate becomes
$\left\lVert\Delta_{\mathbf{W}^{L}}\mathbf{f}\right\rVert_{\text{RMS}}^{2}=\text{tr}(\mathbf{U})\frac{\mathbb{E}\left(\sum_{jj^{\prime}}[\mathbf{Z}^{T}\mathbf{Z}]_{jj^{\prime}}\right)}{\text{tr}(\mathbf{U})}=\mathbb{E}\left(\sum_{jj^{\prime}}[\mathbf{Z}^{T}\mathbf{Z}]_{jj^{\prime}}\right).$ (32)
Similarly for the biases $\mathbf{b}^{L}\in\mathbb{R}^{K}$ we have
$Z_{i}^{(\mathbf{b}^{L})}=\Delta b^{L}_{i}\frac{1}{\sqrt{NK}}\sum_{n}\omega_{ni}$ (33)
and so $Z_{i}^{(\mathbf{b}^{L})}$ are independent for different $i$, and so we can assume the covariance matrix is diagonal. If we assume a diagonal covariance structure over $Z_{i}^{(\mathbf{b}^{L})}$, i.e. $\mathbf{\Sigma}\coloneqq\text{diag}(\boldsymbol{\sigma}^{2})$, then
$||\Delta_{\ell}\mathbf{f}||^{2}_{\text{RMS}}=\sum_{ii^{\prime}}\Sigma_{ii^{\prime}}=\sum_{i=1}^{n}\sigma^{2}_{i}=\sum_{i=1}^{K}\mathbb{E}\left(z_{i}^{(\mathbf{b}^{L})}\right)^{2}=\mathbb{E}\left[\sum_{i=1}^{K}\left(z_{i}^{(\mathbf{b}^{L})}\right)^{2}\right].$ (34)
One might ask at this point how this differs from an IID assumption, i.e. $\mathbf{\Sigma}\coloneqq\sigma^{2}\mathbf{I}$. With an IID assumption, we would have
$||\Delta_{\ell}\mathbf{f}||^{2}_{\text{RMS}}=\sum_{ii^{\prime}}\Sigma_{ii^{\prime}}=n\sigma^{2}=n\mathbb{E}\left(z_{1}^{(\mathbf{b}^{L})}\right)^{2}=n\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{K}\left(z_{i}^{(\mathbf{b}^{L})}\right)^{2}\right]=\mathbb{E}\left[\sum_{i=1}^{K}\left(z_{i}^{(\mathbf{b}^{L})}\right)^{2}\right]$ (35)
which is exactly the same! It turns out that, because we only care about estimating the sum over the variances, assuming that $Z_{i}^{(\mathbf{b}^{L})}$ are IID or just independent are equivalent. Intuitively, consider the difference between making “no assumption” on the covariance matrix versus a diagonal assumption, if we care about the sum over all covariance elements. The diagonal assumption allows us to remove cross-terms from our estimate, since we know their true values are all zero, and hence we remove the variance from the randomness in those cross-terms, giving us a lower variance estimator. But if we consider the remaining positive sum over the variance (diagonal) terms, knowing they’re all equal (an IID assumption) doesn’t provide us with any more useful information, because it doesn’t tell us any more about the value of the total sum, and it doesn’t tell us any more about which values we should pay more attention to (in fact, it tells us to treat them all equally, which is what we already did when we only knew that the covariance was diagonal). If we had more specific information, such as ”the last element’s variance forms 90% of the total variance” then we could use more sophisticated weighted averages to get a lower variance estimate.
In Figure 26, we repeat the width transfer experiments from the main paper but using the tricks we just described, and find that some of the instability / noise in the ResMLP FLeRM plot is now gone.
Figure 26: Width transfer experiments as in main paper, but using output independence tricks given in Section D